A cylindrical drum 2-ft diameter and 4ft height is filled with fluid whose density is 40lb/ft^3.

Determine

(A) the total volume of

(B) its total mass in pounds and slugs

(C) its specific volume and

(D) its specific weight where g=31.90 fps^2


1
Expert's answer
2022-02-10T15:06:41-0500

(1) "V=\pi r^2h"

"V=3.14\times1^2\times4=12.56ft^3"

(2) mass

"M=\rho\times V"

"M=40\times12.56=502.4lb"

1lb=0.031slung


"M=502.4\times0.031=15.5744slug"

(C) specific volume

"v=\frac{1}{\rho}=\frac{1}{40}=0.025ft^3/lb"

(D)

Specific weight

"W=\rho\times g"

"W=40\times31.9=1276lb/ft.s^2"